For what value of a and b will
f(x)={█((x^2-4)/(x-2); x≤2@〖ax〗^2-bx+3; 2≤x≤3@2x-a+b; x≥3)┤
Be continuous at R
1
Expert's answer
2020-05-07T18:08:28-0400

Given function is continuous on R.

So, function f(x) is continuous at x=2 and x=3 also.

So, f(2-) = f(2+) and f(3-)=f(3+).

Now by continuity of f(x) at x=2, we get

"\lim_{x \to 2^-} f(x) = \lim_{x \to 2} \frac{x^2-4}{x-2} = \lim_{x \to 2} (x+2) = 4. \\
\lim_{x \to 2^+} f(x) = (2a)^2-2b+3.\\
\implies 4a^2-2b+3 = 4 \\
\implies 4a^2-2b -1= 0."

Similarly by continuity of f(x) at x=3, we have

"f(3^-)=f(3^+)\\
\implies
9a^2-3b+3=6-a+b \\
\implies 9a^2+a-4b-3=0."

So we get "a^2+a-1=0.
\implies a=\frac{-1\pm\sqrt{5}}{2}."

And

"b = (4a^2-1)\/2 = (4-4a-1)\/2 = (3-4a)\/2."